In a large resturanat an average of 2 out of every 5 customers ask for chilies with their meals. A random sample of 15 customers is selected. Find the probability that i. Exactly 7 customers ask for chilies with their meal ii. Less than 3 customers ask for chilies with their meals. 



1
Expert's answer
2021-11-23T17:55:32-0500

"p = \frac{2}{5} = 0.4 \\

q = 1 - 0.4 = 0.6"

X~Binomial(15,0.4)

"P(X) = C^{15}_x(0.4)^x(0.6)^{15-x}"

i. Exactly 7 customers ask for chilies with their meal

"P(X=7) = C^{15}_7 \times (0.4)^7 \times (0.6)^{15-7} = 0.1770"

ii. Less than 3 customers ask for chilies with their meals.

"P(X<3) = P(X=0) +P(X+1) +P(X=2) \\

P(X=0) = C^{15}_0 \times (0.4)^0 \times (0.6)^{15-0} = 0.00047 \\

P(X=1) = C^{15}_1 \times (0.4)^1 \times (0.6)^{15-1} = 0.00470 \\

P(X=2) = C^{15}_2 \times (0.4)^2 \times (0.6)^{15-2} = 0.02194 \\

P(X<3) = 0.00047 + 0.00470 + 0.02194 = 0.02711"